OpenStudy (anonymous):

The sum of first n terms of the sequence (6, 36, 216, ..., 6n, ...) is 55,986. Under these conditions, whereas log 2 and log 3 = 0.30 = 0.48, the value of log n is: a) 0,78 b) 1,26 c) 1,56 d) 1,08 Help me guys! Do I have to use the sum formula of arithmetical progression ?

OpenStudy (anonymous):

i guess we need n right?

OpenStudy (anonymous):

yep, i guess we need n to find log n.

OpenStudy (anonymous):

\[55968=\frac{6^n-1}{6-1}\] \[5\times 55968=6^n-1\] \[ 299839=6^n\]

OpenStudy (anonymous):

i made a mistake somewhere, let me think

OpenStudy (anonymous):

I did this: (6 + an)n / 2 = 55986 and then: (6 + an)n = 111972

OpenStudy (anonymous):

but it just doesnt make sense. how do i find an and n at the same time?

OpenStudy (anonymous):

n = 6

OpenStudy (anonymous):

so \[\ln(6)=\ln(3)+\ln(2)\] and we are done

OpenStudy (anonymous):

How did you find n = 6??

OpenStudy (anonymous):

i found the 6 by brute force but we can also use the formula

OpenStudy (anonymous):

\[6\frac{6^n-1}{6-1}=55986\] \[6^n-1=\frac{5}{6}\times 55986\] \[6^n=46656\]

OpenStudy (anonymous):

since n is an integer, we can guess and check. if that is too much work then a calculator will find it via \[n=\frac{\ln(46656)}{\ln(6)}\]

OpenStudy (anonymous):

which formula did you use? i dont think ive learnt it